OpenStudy (anonymous):

solve q-7(5/6)=6(1/2)

OpenStudy (anonymous):

q-7(5/6)=6(1/2) q-7(5/6) + 7(5/6) = 6(1/2) + 7(5/6) add 7(5/6) to both sides then simplify

OpenStudy (anonymous):

q=6(1/2) + 7(5/6)

OpenStudy (anonymous):

yep... now simplify....